OpenStudy (anonymous):

vectors: solve x for AB=CD (vectors) A(x,1),B(4,x+3),C(x,x+2) and D(2x,x+6) where should I start?

OpenStudy (kinggeorge):

Is this is saying that A=<x,1>?

OpenStudy (anonymous):

A(x,1) is a point. I guess is AB is (4-x,x+3-1), but it didn't work for me.

OpenStudy (amistre64):

\[\vec{AB}=\vec{CD}\]

OpenStudy (anonymous):

the says the that x=2.

OpenStudy (amistre64):

in other words: B-A = D-C

OpenStudy (amistre64):

of course i see no reason why A-B = C-D would not work equally as well

OpenStudy (anonymous):

yes \[\frac{ \rightarrow }{ AB}=\frac{ \rightarrow }{ CD}\]

OpenStudy (amistre64):

A( x, 1 ) C( x , x+2) -B(-4,-x-3) -D(-2x,-x-6) --------- ----------- (x-4,-2-x) = ( -x , 0x-4) so equate like components x-4 = -x -2-x = -4

OpenStudy (anonymous):

Ok. I did right, but in the end instead put x-4=-x I put x-4 -x =0 <--- wrong. Again thank you amistre64.

OpenStudy (amistre64):

youre welcome